OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method.

OpenStudy (lgbasallote):

intercepts method so find the x and y intercept x -intercept let y= 0 3x + 9(0) = -18 3x = -18 x = -18/3 x = -6 y-intercept let x = 0 3(0) + 9y = -18 9y = -18 y = -18/9 y = -2 therefore x-intercept is (-6,0) y-intercept is (0,-2) plot it and connect the dots

OpenStudy (anonymous):

0+9y = -18 y=-2 x=0 3x=-18 x=-6 y=0